# Earliest Galaxy Evolution in the CANUCS+Technicolor fields: Galaxy Properties at $z\sim 10-16$ seen with the Full NIRCam Medium and Broad Band Filters

### v1

@@ 5.3.2 Rest UV slopes
Previous JWST studies have found rest-UV slopes, $\beta_{\rm UV}$, are blue at the highest redshifts (e.g., F. Cullen et al., 2024; M. W. Topping et al., 2024), and some of them suggest galaxies at $z>10$ are remarkably blue and could be completely dust-free (F. Cullen et al., 2024). On the other hand, galaxies selected from our CANUCS+TEC data and from JOF (B. Robertson et al., 2024) are typically less blue ($\langle\beta_{\rm UV}\rangle=-2.2$ in CANUCS+TEC and $\langle\beta_{\rm UV}\rangle=-2.5$ in JOF; see also N. Martis et al. 2025 for the high-$z$ red galaxy population in CANUCS+TEC fields). This could indicate that rest-UV slope studies can also be biased in some cases with BB-only data due to bluer galaxies being easier to distinguish from low-$z$ interlopers and to fewer filters to compute $\beta_{\rm UV}$.
Figure 7 shows the distribution of the robust high-$z$ sample (CANUCS+TEC, red squares; JOF, black crosses) in the $\beta_{\rm UV}$-$M_{\rm UV}$ plane, compared to the $z\sim 11.5$ galaxy sample from F. Cullen et al. (2024) (light blue circles), which mostly used BB-only photometry data in the sample selection. Interestingly, the high-$z$ galaxies selected with MB+BB data distributes in the relatively less blue and fainter locus in the $\beta_{\rm UV}$-$M_{\rm UV}$ plane ($M_{\rm UV}\gtrsim-19$ mag and $\beta_{\rm UV}\gtrsim-2.5$), and this population is absent in the F. Cullen et al. (2024) sample. This can be because relatively red high-$z$ galaxies close to the detection limit of the observation can be degenerate with the low-$z$ Balmer break galaxy solution, particularly when only BB data is available, and they could be missed from high-$z$ selection with BB-only data. A representative example is the relatively red $z\sim 15.4$ galaxy candidate CANUCS-3214552 (Sec. 3.1), whose low-$z$ probability is substantial ($P_{\rm BB}(z<7)=0.13$) when only BB data is used, but negligible with the full MB+BB data ($P(z<7)=0.005$). Therefore, BB-only selection of high-$z$ galaxy candidates can be relatively incomplete for faint and less blue galaxies.
Figure 7: The $\beta_{\rm UV}$-$M_{\rm UV}$ relation. High-$z$ galaxy candidates based on MB+BB observations (red squares and black crosses) are less blue than those of a comparison sample mostly based on BB-only data (F. Cullen et al., 2024, cyan circles). In particular, there is a population of faint ($M_{\rm UV}\gtrsim-19$), relatively red ($\beta_{\rm UV}\gtrsim-2.5$) galaxies that is largely absent from the F. Cullen et al. (2024) sample. False Positives (blue squares) cover most of the observed range of $\beta_{\rm UV}$. The linear regression of the $\beta_{\rm UV}$-$M_{\rm UV}$ relation based on the main MB+BB selection sample (red dashed line) is less blue than found previously (cyan dashed line).
@@
fixing the slope $d\beta_{\rm UV}/dM_{\rm UV}=-0.15$ (red dashed line in Figure 7). The intercept of $\beta_{\rm UV}(M_{\rm UV}=-19)=-2.36\pm 0.04$ at $\langle z\rangle=11.5$ is redder by $\Delta\beta\sim 0.3$ than F. Cullen et al. (2024, cyan dashed line) at this redshift.
@@ 5.4 Future galaxy surveys at $z\gtrsim 10$
By exploiting the robust high-$z$ candidates selected with rich MB+BB data, we demonstrate the importance of MBs in studies of $z\gtrsim 10$ galaxies. However, the number of known galaxies is still small and enlarging the sample size is necessary. The bright-end at $z\gtrsim 14$ needs particular caveats, but bright $z\gtrsim 14$ galaxy candidates are not always interlopers (e.g., GS-z14-0, S. Carniani et al. 2024; K. N. Hainline et al. 2024; MoM-z14, R. P. Naidu et al. 2025). Securing a high-$z$ bright galaxy sample is crucial in various aspects (e.g., bright-end of LFs, scaling relations such as size-$M_{\rm UV}$ relation), and the effect of CV can be considerable in such a rare population search (Sec. 5.1).
Figure 8: UV LFs at $z\sim 14$ from the two surveys with extensive MB+BB filter observations (CANUCS+TEC, this work; JOF, B. Robertson et al. 2024), compared with the MINERVA survey detectability of $z\sim 14$ galaxies. The two UV LF measurements agree at the faint end ($M_{\rm UV}>-19$) but show a large discrepancy at the bright end. The JWST Cycle 4 MINERVA survey will provide $\sim 500$ arcmin2 with almost all MB+BB filters across four independent line of sights, and will give a tight constraint at the bright end of the highest-$z$ UV LF by building a very robust bright galaxy sample at $z\gtrsim 14$ (blue square area).
[…]
@@ 6 Summary
    We obtain a sample of eight galaxies at $169.5<z<$ 16 with the selection 50 % complete down to $M_{\rm UV}\sim-18.7$ mag over the full redshift range (Figure 2). The highest redshift candidate (CANUCS-3214552) is found at $z_{\rm phot}\sim 15.4$ with a sharp dropout between F210M and F200W (Figure 1). The galaxy CANUCS-3214552 is faint ($M_{\rm UV}=-19$ mag) and relatively red ($\beta=-1.8$), and, due to the relatively red color, it could not be identified as a reliable high-$z$ candidate only with the BB filter data. Also it shows no variability between Cycle 1 and Cycle 2 NIRCam observations, thus CANUCS-3214552 is one of the most reliable candidates of $z>15$ galaxies, which can be secured only with multi-epoch, full NIRCam MB+BB observations (Sec. 3). It is absolutely necessary to follow-up this galaxy with NIRSpec Prism to confirm its redshift and probe the detailed physical properties.
@@
    The UV LF from CANUCS+TEC data shows a moderate redshift evolution at $z>10$. The lack of relatively bright galaxies (brighter than $M_{\rm UV}=-20$ mag) and the paucity of $z>12$ galaxies over $\sim 23\ {\rm arcmin}^{2}$ places tight constraints on the redshift evolution, and the redshift-dependent UV LF normalization parameter is found to be $\log\phi_{\star}(z)=-4.89-0.21(z-10)$ (Table 3 and Figure 3). Our UV LF at $1610<z<$ 16 is $\sim 0.6$ dex lower than a previous work in the JADES Origins Field (JOF), where similarly rich NIRCam MB+BB filters are available, suggesting that the cosmic variance can affect galaxy abundance measurement by at least 0.6 dex in the very high-$z$ universe (Sec. 4.1, 4.2).
@@
    The log-linear rate of the $\rho_{\rm UV}$ redshift evolution at $z\sim 10-16$ from this work is $-0.21$. This is somewhat shallower than at lower redshift of $z\sim 8-10$ and is consistent with the previous work at the similar redshift in the JOF. Comparing with a prediction of $\rho_{\rm UV}$ evolution assuming the constant star formation efficiency (SFE), it could suggest the SFE starts to deviate from the simple constant extrapolation at $z\gtrsim 11$ (Figure 4 and Sec. 4.3).
@@
    By comparing high-$z$ galaxy candidates from the BB-only degraded catalog and from the full MB+BB catalog, the low-$z$ interloper contamination is found to be non-negligible when selecting $z>9.5$ galaxies based only on the BB data. The major population of interlopers is very dusty EELGs that starts to populate at intermediate redshift of $z>2$, and roughly a half of $z>9.5$ galaxy candidates from BB-only photometry could be low-$z$ interlopers (Sec. 5). The most clean galaxy samples at $z\gtrsim 10$ identified from the rich MB+BB observations suggest that highest-$z$ galaxies could be rarer, more compact, and less blue than previously claimed: the UV LF normalization could be overestimated by $\sim 0.6$ dex when BB-only data is used (Figure 5); the low-$z$ interlopers appears systematically larger than real high-$z$ sources, leading to an overestimation of the size-$M_{\rm UV}$ relation by $\sim 0.6$ dex (Figure 6); the main high-$z$ galaxy sample from the MB+BB selection is less blue while low-$z$ interlopers can masquerade very blue color in the BB-only SEDs, resulting in a systematically bluer estimation of the $\beta_{\rm UV}-M_{\rm UV}$ relation by $\Delta\beta\sim-0.3$ (Figure 7).
@@
This work is based on observations made with the NASA/ESA/CSA JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-03127 for JWST. This research was supported by grants 18JWST-GTO1, 23JWGO2A13, and 23JWGO2B15 from the Canadian Space Agency (CSA), and funding from the Natural Sciences and Engineering Research Council of Canada (NSERC). Support for program JWST-GO-03362 was provided through a grant from the STScI under NASA contract NAS5-03127. This research used the Canadian Advanced Network For Astronomy Research (CANFAR) operated in partnership by the Canadian Astronomy Data Centre and The Digital Research Alliance of Canada with support from the National Research Council of Canada, the Canadian Space Agency, CANARIE and the Canadian Foundation for Innovation. YA is supported by JSPS KAKENHI Grant Number 23H00131. MB, NM, GF, JJ acknowledge support from the ERC Grant FIRSTLIGHT, Slovenian national research agency ARIS through grants N1-0238 and P1-0188, and the program HST-GO-16667, provided through a grant from the STScI under NASA contract NAS5-26555. D.M. and A.L.R. acknowledge support from program JWST-GO-03362, provided through a grant from the STScI under NASA contract NAS5-03127.
@@ A.1 CANUCS-2217931
Figure 9: A False Positive CANUCS-2217931. Top: cutouts of the source in all NIRCam images. BB images are displayed with blue frames, while MB images are with red frames. Cutouts are $1.\!\!^{\prime\prime}6$ on the side. Bottom left: SED of the source. Photometry in BB and MB filters are shown by blue and red open squares, respectively. The best-fit model spectrum when only BB data are used is shown by blue slid curve, while the best model for all MB+BB data is shown by black. Bottom right: $P(z)$ of the source. Blue curve shows the $P(z)$ when only BB data is used, and red shows that for the all MB+BB data SED.
@@ A.2 CANUCS-2225353
CANUCS-2225353 is a $z\sim 16$ galaxy candidate found in the BB-only catalog with almost zero low-$z$ solution probability ($8.1\times 10^{-5}$). The SED shows a clear break in the F200W-F277W color and apparently blue SED from F277W to F [^4] W, which favors $z\sim 16$ solutions when only BB data is used (blue solid curves in Figure 10). With the MB observations, though, the SED shows extreme excess in F300M and F360M photometry, which clearly indicates the source is a very dusty EELG at $z\sim 5$. The source is detected in F150W with S/N $=6$, thus CANUCS-2225353 is removed from the high-$z$ sample of the BB-only catalog in Section 5.2.

### v2

@@ 5.3.2 Rest UV slopes
Previous JWST studies have found rest-UV slopes, $\beta_{\rm UV}$, are blue at the highest redshifts (e.g., Cullen2024MNRAS; Topping2024), and some of them suggest galaxies at $z>10$ are remarkably blue and could be completely dust-free (Cullen2024MNRAS). On the other hand, galaxies selected from our CANUCS+TEC data and from JOF (Robertson2024ApJ) are typically less blue ($\langle\beta_{\rm UV}\rangle=-2.2$ in CANUCS+TEC and $\langle\beta_{\rm UV}\rangle=-2.5$ in JOF; see also Martis2025arXiv for the high-$z$ red galaxy population in CANUCS+TEC fields). This could indicate that rest-UV slope studies can also be biased in some cases with BB-only data due to bluer galaxies being easier to distinguish from low-$z$ interlopers and to fewer filters to compute $\beta_{\rm UV}$.
Figure 8 shows the distribution of the robust high-$z$ sample (CANUCS+TEC, red squares; JOF, black crosses) in the $\beta_{\rm UV}$-$M_{\rm UV}$ plane, compared to the $z\sim 11.5$ galaxy sample from Cullen2024MNRAS (light blue circles), which mostly used BB-only photometry data in the sample selection. Interestingly, the high-$z$ galaxies selected with MB+BB data distributes in the relatively less blue and fainter locus in the $\beta_{\rm UV}$-$M_{\rm UV}$ plane ($M_{\rm UV}\gtrsim-19$ mag and $\beta_{\rm UV}\gtrsim-2.5$), and this population is absent in the Cullen2024MNRAS sample. This can be because relatively red high-$z$ galaxies close to the detection limit of the observation can be degenerate with the low-$z$ Balmer break galaxy solution, particularly when only BB data is available, and they could be missed from high-$z$ selection with BB-only data. A representative example is the relatively red $z\sim 15.4$ galaxy candidate CANUCS-3214552 (Sec. 3.1), whose low-$z$ probability is substantial ($P_{\rm BB}(z<7)=0.13$) when only BB data is used, but negligible with the full MB+BB data ($P(z<7)=0.005$). Therefore, BB-only selection of high-$z$ galaxy candidates can be relatively incomplete for faint and less blue galaxies.
Figure 8: The $\beta_{\rm UV}$-$M_{\rm UV}$ relation. High-$z$ galaxy candidates based on MB+BB observations (red squares and black crosses) are less blue than those of a comparison sample mostly based on BB-only data (Cullen2024MNRAS, cyan circles). In particular, there is a population of faint ($M_{\rm UV}\gtrsim-19$), relatively red ($\beta_{\rm UV}\gtrsim-2.5$) galaxies that is largely absent from the Cullen2024MNRAS sample. False Positives (blue squares) cover most of the observed range of $\beta_{\rm UV}$. The linear regression of the $\beta_{\rm UV}$-$M_{\rm UV}$ relation based on the main MB+BB selection sample (red dashed line) is less blue than found previously (cyan dashed line).
@@
fixing the slope $d\beta_{\rm UV}/dM_{\rm UV}=-0.15$ (red dashed line in Figure 8). The intercept of $\beta_{\rm UV}(M_{\rm UV}=-19)=-2.36\pm 0.04$ at $\langle z\rangle=11.5$ is redder by $\Delta\beta\sim 0.3$ than Cullen2024MNRAS at this redshift.
@@ 5.4 Future galaxy surveys at $z\gtrsim 10$
By exploiting the robust high-$z$ candidates selected with rich MB+BB data, we demonstrate the importance of MBs in studies of $z\gtrsim 10$ galaxies. However, the number of known galaxies is still small and enlarging the sample size is necessary. The bright-end at $z\gtrsim 14$ needs particular caveats, but bright $z\gtrsim 14$ galaxy candidates are not always interlopers (e.g., GS-z14-0, Carniani2024; Hainline2024; MoM-z14, Naidu2025). Securing a high-$z$ bright galaxy sample is crucial in various aspects (e.g., bright-end of LFs, scaling relations such as size-$M_{\rm UV}$ relation), and the effect of CV can be considerable in such a rare population search (Sec. 5.1).
Figure 9: UV LFs at $z\sim 14$ from the two surveys with extensive MB+BB filter observations (CANUCS+TEC, this work; JOF, Robertson2024ApJ), compared with the MINERVA survey detectability of $z\sim 14$ galaxies. The two UV LF measurements agree at the faint end ($M_{\rm UV}>-19$) but show a large discrepancy at the bright end. The JWST Cycle 4 MINERVA survey will provide $\sim 500$ arcmin2 with almost all MB+BB filters across four independent line of sights, and will give a tight constraint at the bright end of the highest-$z$ UV LF by building a very robust bright galaxy sample at $z\gtrsim 14$ (blue square area).
Figure 9 compares the UV LF measurements at $z\sim 14$ from this work (black) and from JOF (green; Robertson2024ApJ), both are based on MB+BB observations. Although they agree with each other at the faint-end, the bright-end of UV LFs is poorly constrained due to their paucity and the two measurements are largely discrepant at $M_{\rm UV}\lesssim-20$ mag. The large discrepant at the bright-end is mainly due to the presence of one bright galaxy, GS-z14-0, in the small footprint of JOF. The galaxy is thus far the brightest galaxy at this redshift range even across the all JADES field-of-views that covers $\sim\times 15$ area of the JOF, and it can significantly affect the number density of bright galaxies in JOF. It is thus required to search for these bright high-$z$ galaxies with MB+BBs over wide areas in independent lines of sight, mitigating the effect of field-to-field variance.
@@ 6 Summary
    We obtain a sample of eight galaxies at $9.5<z<16$ with the selection 50 % complete down to $M_{\rm UV}\sim-18.7$ mag over the full redshift range (Figure 3). The UV LF from CANUCS+TEC data shows a moderate redshift evolution at $z>10$. The lack of relatively bright galaxies (brighter than $M_{\rm UV}=-20$ mag) and the paucity of $z>12$ galaxies over $\sim 23\ {\rm arcmin}^{2}$ places tight constraints on the redshift evolution, and the redshift-dependent UV LF normalization parameter is found to be $\log\phi_{\star}(z)=-4.89-0.21(z-10)$ (Table 3 and Figure 4). Our UV LF at $10<z<16$ is $\sim 0.6$ dex lower than a previous work in the JADES Origins Field (JOF), where similarly rich NIRCam MB+BB filters are available, suggesting that the cosmic variance can affect galaxy abundance measurement by at least 0.6 dex in the very high-$z$ universe (Sec. 4.1, 4.2).
@@
    The highest redshift candidate in our sample (CANUCS-3214552) is found at $z_{\rm phot}\sim 15.4$ with a sharp dropout between F210M and F200W (Figure 2). The galaxy CANUCS-3214552 is faint ($M_{\rm UV}=-19$ mag) and relatively red ($\beta=-1.8$), and, due to the relatively red color, it could not be identified as a reliable high-$z$ candidate only with the BB filter data. Also it shows no variability between Cycle 1 and Cycle 2 NIRCam observations, thus CANUCS-3214552 is one of the most reliable candidates of $z>15$ galaxies, which can be secured only with multi-epoch, full NIRCam MB+BB observations (Sec. 3). It is absolutely necessary to follow-up this galaxy with NIRSpec Prism to confirm its redshift and probe the detailed physical properties.
@@
    The log-linear rate of the $\rho_{\rm UV}$ redshift evolution at $z\sim 10-16$ from this work is 0.21-0.21. This is somewhat shallower than at lower redshift of $z\sim 8-10$ and is consistent with the previous work at the similar redshift in the JOF. Comparing with a prediction of $\rho_{\rm UV}$ evolution assuming the constant star formation efficiency (SFE), it could suggest the SFE starts to deviate from the simple constant extrapolation at $z\gtrsim 11$ (Figure 5 and Sec. 4.3).
@@
    By comparing high-$z$ galaxy candidates from the BB-only degraded catalog and from the full MB+BB catalog, the low-$z$ interloper contamination is found to be non-negligible when selecting $z>9.5$ galaxies based only on the BB data. The major population of interlopers is very dusty EELGs that starts to populate at intermediate redshift of $z>2$, and roughly a half of $z>9.5$ galaxy candidates from BB-only photometry could be low-$z$ interlopers (Sec. 5).
@@
    The most clean galaxy samples at $z\gtrsim 10$ identified from the rich MB+BB observations suggest that highest-$z$ galaxies could be rarer, more compact, and less blue than previously claimed: the UV LF normalization could be overestimated by $\sim 0.6$ dex when BB-only data is used (Figure 6); the low-$z$ interlopers appears systematically larger than real high-$z$ sources, leading to an overestimation of the size-$M_{\rm UV}$ relation by $\sim 0.6$ dex (Figure 7); the main high-$z$ galaxy sample from the MB+BB selection is less blue while low-$z$ interlopers can masquerade very blue color in the BB-only SEDs, resulting in a systematically bluer estimation of the $\beta_{\rm UV}-M_{\rm UV}$ relation by $\Delta\beta\sim-0.3$ (Figure 8).
  7. 7.
@@
We thank the anonymous referees for their thoughtful comments and suggestions to improve the quality of the paper. This work is based on observations made with the NASA/ESA/CSA JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-03127 for JWST. This research was supported by grants 18JWST-GTO1, 23JWGO2A13, and 23JWGO2B15 from the Canadian Space Agency (CSA), and funding from the Natural Sciences and Engineering Research Council of Canada (NSERC). Support for program JWST-GO-03362 was provided through a grant from the STScI under NASA contract NAS5-03127. This research used the Canadian Advanced Network For Astronomy Research (CANFAR) operated in partnership by the Canadian Astronomy Data Centre and The Digital Research Alliance of Canada with support from the National Research Council of Canada, the Canadian Space Agency, CANARIE and the Canadian Foundation for Innovation. The Dunlap Institute is funded through an endowment established by the David Dunlap family and the University of Toronto. YA is supported by JSPS KAKENHI Grant Number 23H00131. MB, NM, GF, $J$ acknowledge support from the ERC Grant FIRSTLIGHT, Slovenian national research agency ARIS through grants N1-0238 and P1-0188, and the program HST-GO-16667, provided through a grant from the STScI under NASA contract NAS5-26555. D.M. and A.L.R. acknowledge support from program JWST-GO-03362, provided through a grant from the STScI under NASA contract NAS5-03127.
Data Availability
The data products used in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute, via https://doi.org/10.17909/ph4n-6n76 (catalog doi: 10.17909/ph4n-6n76) and https://doi.org/10.17909/cyh7-mm53 (catalog doi: 10.17909/cyh7-mm53). The CANUCS data release HLSP DOI is https://doi.org/10.17909/18nv-np70 (catalog doi:10.17909/18nv-np70). The DR1 data release products can also be found at https://niriss.github.io/data_release1.html (catalog the CANUCS website).
@@ A.1 CANUCS-2217931
Figure 10: A False Positive CANUCS-2217931. Top: cutouts of the source in all NIRCam images. $B$ images are displayed with blue frames, while MB images are with red frames. Cutouts are $1.\!\!^{\prime\prime}6$ on the side. Bottom left: SED of the source. Photometry in BB and MB filters are shown by blue and red open squares, respectively. The best-fit model spectrum when only BB data are used is shown by blue slid curve, while the best model for all MB+BB data is shown by black. Bottom right: $P(z)$ of the source. Blue curve shows the $P(z)$ when only BB data is used, and red shows that for the all MB+BB data SED.
@@ A.2 CANUCS-2225353
CANUCS-2225353 is a $z\sim 16$ galaxy candidate found in the BB-only catalog with almost zero low-$z$ solution probability ($8.1\times 10^{-5}$). The SED shows a clear break in the F200W-F277W color and apparently blue SED from F277W to F [^4] W, which favors $z\sim 16$ solutions when only BB data is used (blue solid curves in Figure 11). With the MB observations, though, the SED shows extreme excess in F300M and F360M photometry, which clearly indicates the source is a very dusty EELG at $z\sim 5$. The source is detected in F150W with S/N =6=6, thus CANUCS-2225353 is removed from the high-$z$ sample of the BB-only catalog in Section 5.2.